OpenStudy (anonymous):

Find an equation in standard form for the ellipse with the vertical major axis of length 18, and minor axis of length 16.

OpenStudy (anonymous):

\[\frac{ y-a }{ 9^2 }+\frac{ x-b }{ 4^2 }=1\] where (b,a) is the point which is the centre of the ellipse and.